• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nilanga

    Log entry time 16:13:59 on April22,2001

    Entry number 61044

    keyword=Shift Summary

    8.00: Continuing with the hyper-white scan
    Theta 43.58
    P = 1.0228
    7.44: LH2 3783
    8.26: optics 3784

    P= 1.0023
    8.44: LH2 3785
    9.18: optics 3786

    P=0.9823
    9.32: optics 3787
    9.43: LH2 3788
    *********
    L-arm Q1 field was about 1.10-3: John LeRose suggested to adjust the set current
    by hand to get the desired value: this this.

    Asked John whether it is usefull to take open collimator runs in addition to 6msr runs. He thought it was a good idea and we will do this from now on.

    **********
    10.07: LH2 open collimator 3789

    Theta= 46.4


    *********
    y_tg resolution for L-Arm factort of 2 worse than R-arm.

    Use single foil 12C target to check
    ********

    p=1.065 GeV (#11)
    10.52: single foil 12C: 3790
    11.06: 9-foil optics: 3791
    11.17: LH2: open collimator: 3792
    11.43: LH2: 6msr: 3793

    P= 1.0437

    12.23: LH2 6msr: 3794
    12.56: LH2 open: 3795

    P=1.0228
    13.24: LH2 open: 3796
    13.45: LH2 6msr: 3797
    14.15: optics : 3798
    14.22: optics, beam spot=0,0: 3799
    14.29: Optics, beam spot=+2,0: 3800
    14.44: LH2: 6msr: 3801



    ********************

    L-Arm resolution problem still not solved:

    took an access to check whether there is anything blocking the spectrometer:

    The entrance window shield on L-Arm is still there.

    Discussed with Joe whether we should remove it now: We should not change things in the middle of the run: Wait till the end of the run

    ********************